how many trios can be selected from a group of 5 musicians?
Answer by jim_thompson5910(35256)

Assuming order doesn't matter, then...

n C r = (n!)/(r!(n-r)!)

5 C 3 = (5!)/(3!*(5-3)!)

5 C 3 = (5!)/(3!*2!)

5 C 3 = (5*4*3!)/(3!*2!)

5 C 3 = (5*4)/(2!)

5 C 3 = (5*4)/(2*1)

5 C 3 = (20)/(2)

5 C 3 = 10

Assuming order doesn't matter, then there are 10 different ways to do this.